### v1

# Boosted second moment method in random regular graphs

Abstract.
Determining the asymptotic independence ratio of random regular graphs is a key challenge in the area of sparse random graphs. Due to the interpolation method, very good upper bounds are known, which are actually known to be sharp for sufficiently large degrees.
This paper provides explicit lower bounds for given degrees $d$, beating the previous best bounds for $d\geq 10$. The starting point is a second moment argument that we can boost by arguing that the obtained independent set has a certain spatial Markov property. One can then exploit this property by making local modifications to the independent set, resulting in substantial improvements.
Our approach can also be used to prove the existence of other objects in random regular graphs. To demonstrate this, we consider the problem of decomposing random regular graphs into stars.
The first author was supported by the National Research, Development and Innovation Office (NRDI; grant KKP 137490).
The second author was supported by the MTA-Rényi Counting in Sparse Graphs “Momentum” Research Group, by NRDI (grant KKP 138270), and by the Hungarian Academy of Sciences (János Bolyai Scholarship).
1. Introduction
This paper is concerned with random regular graphs: let $\mathbb{G}_{N,d}$ be a uniform random graph among all $d$-regular simple graphs on the vertex set $\{1,\ldots,N\}$. A standard Azuma-type argument shows that the independence ratio of $\mathbb{G}_{N,d}$ is highly concentrated around its expectation for any fixed $d$ and $N$. A classic result [4] shows that for each degree $d\geq 3$, there exists a constant $\alpha^{\star}_{d}$ such that the independence ratio of the random $d$-regular graph $\mathbb{G}_{N,d}$ converges in probability to $\alpha^{\star}_{d}$ as $N\to\infty$.
Determining $\alpha^{\star}_{d}$, or at least providing good bounds for it, is a major challenge of the area. In [11, 10] Frieze and Łuczak used a second moment argument to give accurate asymptotics for $\alpha^{\star}_{d}$ as $d\to\infty$:
(1) $\alpha^{\star}_{d}=\frac{2}{d}\bigg(\log d-\log\log d+1-\log 2+o_{d}(1)\bigg).$
Non-rigorous statistical physics methods suggested that the precise value of $\alpha^{\star}_{d}$ is given by a certain 1-RSB (1-step replica symmetry breaking) formula for $d\geq 20$ [3]. In their seminal work [8] Ding, Sun, and Sly confirmed this prediction for very large $d$. Below $20$ there is a phase transition: for $d\leq 19$ one can prove upper bounds that are stronger than the 1-RSB formula [12]. These RSB bounds are likely to be extremely close to the true value.
As for lower bounds for specific degrees, however, there is still a long way to go to close the gaps. Most results in this direction attained lower bounds by analyzing certain (local) processes/algorithms. [^1] Many results actually apply to the more general setting of graphs of (essentially) large girth. A noteable tool is Wormald’s differential equation method. [^2] This method was introduced in [15]. See [14] for a concise treatment. However, when the algorithms are more sophisticated, then it tends to be very challenging to analyze them, often requiring considerable computer assistance. Due to such difficulties, some approaches treat only the smallest cases $d=3$ or $d=3,4$. Currently, the best lower bounds are due to Csóka [5] for $d=3,4$; to Duckworth–Zito [9] for $d=5,6,7$; and to Wormald [15] for $d\geq 8$, apart from $d=16$ for which degree a better bound was achieved via star decompositions in [6].
In this paper we revisit the second moment method. As we mentioned, the asymptotic formula (1) for $\alpha^{\star}_{d}$ was attained using this method [10]. In fact, that result built on an earlier work [11] regarding the independence ratio of sparse Erdős–Rényi random graphs, and the second moment method was actually used in that earlier work. Our approach applies the second moment method directly to random regular graphs. Instead of an asymptotic formula, this gives us explicit lower bounds on $\alpha^{\star}_{d}$ for specific degrees $d$. These bounds would already provide new results, but we can boost them further with the following additional step. One can show that the second moment method actually proves the existence of independent sets with a certain spatial Markov property. This allows us to precisely calculate how much improvement we may obtain by making local modifications to our independent set; see Section 1.2 for details. Table 1 shows our new bounds compared to the best bounds previously known. We also included the 1-RSB upper bounds in the table (which are conjectured to be sharp for $d\geq 20$). Our lower bounds are fairly close: for $d=500$ we have $0.0190$ (lower) versus $0.0193$ (upper).
degree best previous low.bnd our low.bnd best upp.bnd
$10$ [15]$0.2573$ $0.258371$ $0.281105$
$11$ [15]$0.2447$ $0.246775$ $0.268311$
$12$ [15]$0.2335$ $0.236447$ $0.256856$
$13$ [15]$0.2234$ $0.227167$ $0.246525$
$14$ [15]$0.2143$ $0.218767$ $0.237148$
$15$ [15]$0.2061$ $0.211113$ $0.228588$
$16$ [6]$0.2000$ $0.204101$ $0.220737$
$17$ [15]$0.1916$ $0.197645$ $0.213502$
$18$ [15]$0.1852$ $0.191674$ $0.206810$
1919 [15]$0.1793$ $0.186131$ $0.200597$
2020 [15]$0.1738$ $0.180967$ $0.194809$
$50$ [15]$0.0951$ $0.104315$ $0.109797$
$100$ [15]$0.0572$ $0.065030$ $0.067446$
$200$ [15, Thm 4]$0.0260$ $0.039010$ $0.040018$
$500$ [15, Thm 4]$0.0123$ $0.019005$ $0.019307$
$1000$ [15, Thm 4]$0.0068$ $0.010771$ $0.010890$
$2000$ [15, Thm 4]$0.0037$ $0.006014$ $0.006061$
$5000$ [15, Thm 4]$0.0017$ $0.002736$ $0.002750$
$10000$ [15, Thm 4]$0.0009$ $0.001493$ $0.001498$
Table 1. Comparing bounds on the independence ratio $\alpha^{\star}_{d}$
We point out that most papers in the literature only reported their bounds for a small number of degrees. In fact, some of these papers were actually able to carry out the required computer-assisted analysis only for small degrees. For good explicit bounds for degrees $d>20$, one has had to use a weaker bound provided by an explicit formula given in [15, Theorem 4].
Although our bounds do not come in the shape of a closed formula, it is relatively simple to compute the exact values with a computer. Our short SageMath code is able to quickly compute the bound for any given degree $d\leq 500,000$. We uploaded the code to the following GitHub repository.
The reader may run our code on the following interactive website, which also contains a table of our bounds for degrees up to $200$.
In the remainder of the introduction, we give a brief overview of our approach.
1.1. Labelings and entropy
Suppose that $\mathcal{X}$ is a finite set of labels. Given a $d$-regular graph $G$, an $\mathcal{X}$-labeling of $G$ is a mapping $\ell\colon V(G)\to\mathcal{X}$. If $G$ is finite, one can consider the local statistics of a labeling. For instance, we may define the edge statistics of $\ell$ as the distribution $\pi$ of the pair $\big(\ell(u),\ell(v)\big)$, where $uv$ is a uniform random (directed) edge of $G$. Note that $\pi$ is an exchangeable distribution on $\mathcal{X}\times\mathcal{X}$ (i.e., it is invariant under exchanging the two coordinates). In particular, the two marginals of $\pi$ coincide; let us denote this marginal distribution by $\pi_{\mathrm{vtx}}$. (Since $G$ is regular, $\pi_{\mathrm{vtx}}$ is simply the distribution of $\ell(v)$ for a uniform random vertex $v$.) We denote the probabilities defining $\pi$ and $\pi_{\mathrm{vtx}}$ as follows:
$p_{x,y}$ $=p_{y,x}=\pi\big(\{(x,y)\}\big);$
$p_{x}$ $=\sum_{y\in\mathcal{X}}p_{x,y}=\pi_{\mathrm{vtx}}\big(\{x\}\big).$
Later on we will need the Shannon entropy of $\pi$ and $\pi_{\mathrm{vtx}}$ defined as follows:
(2) $H(\pi)=\sum_{x,y\in\mathcal{X}}h\big(p_{x,y}\big)\quad\text{and}\quad H(\pi_{\mathrm{vtx}})=\sum_{x\in\mathcal{X}}h\big(p_{x}\big),\\ \text{where }h(x)=\begin{cases}-x\log x&\text{if }x\in(0,1];\\ 0&\text{if }x=0.\end{cases}$
Many fundamental problems in graph theory can be phrased as questions about whether a graph has a labeling with certain edge statistics. For instance, the independence ratio of $G$ is at least $\alpha$ if and only if $G$ has $\{0,1\}$-labeling with edge statistics
$p_{1,1}=0\quad\text{and}\quad p_{0,1}=p_{1,0}\geq\alpha.$
Recall that $\mathbb{G}_{N,d}$ denotes the uniform random graph among all $d$-regular simple graphs on the vertex set $\{1,\ldots,N\}$. This is closely related to the random multigraph $\mathbb{G}^{\textrm{conf}}_{N,d}$ produced by the so-called configuration model. The exponential rate of the expected number of labelings of $\mathbb{G}^{\textrm{conf}}_{N,d}$ with edge statistics close to a given $\pi$ can be seen to be equal to
$\Sigma(\pi)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\frac{d}{2}H(\pi)-(d-1)H(\pi_{\mathrm{vtx}}).$
A second moment argument (supplemented by an Azuma-type concentration result) gives the following result.
Claim.
Let $\mathcal{P}_{\mathrm{edge}}^{\mathcal{X}}$ denote the set of exchangeable distributions on $\mathcal{X}\times\mathcal{X}$. We say that $\pi\in\mathcal{P}_{\mathrm{edge}}^{\mathcal{X}}$ satisfies the second moment condition if
[…]
If $\pi\in\mathcal{P}_{\mathrm{edge}}^{\mathcal{X}}$ satisfies the second moment condition (3), then, for any $\varepsilon>0$, $\mathbb{G}_{N,d}$ a.a.s. has an $\mathcal{X}$-labeling with edge statistics $\varepsilon$-close to $\pi$.
Next we explain how one can boost this result by considering larger neighborhood statistics. We start with defining the notion of tree-indexed Markov chains. Let $T_{d}$ denote the $d$-regular infinite tree. A $T_{d}$-indexed Markov chain is a random labeling $X\colon V(T_{d})\to\mathcal{X}$ with the property that, conditioned on the label $X_{v}$ of a vertex $v\in V(T_{d})$, $X$ is conditionally independent on the $d$ subtrees obtained after deleting $v$. (Similar properties are often referred to as spatial Markov properties, while the authors of [1] call such random labelings vertex-Markov processes.) It is easy to see that the distribution $\mu$ of $X$ is uniquely determined by its marginal on two neighboring vertices, which can be considered as some $\pi\in\mathcal{P}_{\mathrm{edge}}^{\mathcal{X}}$. For $R\in\mathbb{N}$, let $\mu_{R}$ denote the marginal of $\mu$ corresponding to an $R$-ball of $T_{d}$. It is not hard to show that the following stronger version of the second moment method also holds true; see [1, Thm 1.15].
Claim.
Let $\mu$ be a $T_{d}$-indexed Markov chain with edge marginal $\pi\in\mathcal{P}_{\mathrm{edge}}^{\mathcal{X}}$. If $\pi$ satisfies the second moment condition (3), then, for any $\varepsilon>0$ and any $R\in\mathbb{N}$, $\mathbb{G}_{N,d}$ a.a.s. has an $\mathcal{X}$-labeling with $R$-neighborhood statistics $\varepsilon$-close to $\mu_{R}$.
In essence, the second moment condition guarantees the existence of a labeling that satisfies the spatial Markov property in the sense of local statistics. Therefore, we know what our labeling looks like in larger neighborhoods. Then we can use (randomized) local algorithms to produce new labelings: the new label of a vertex $v$ is computed based on the original labels in a neighborhood of $v$. The point is that we can compute the local statistics of the new labeling using the large-neighborhood statistics of the original one. Next we demonstrate this approach in the context of independent sets.
1.2. Independent sets
We already pointed out that an independent set can be thought of as a $\{0,1\}$-labeling with no adjacent $1$’s. So in the edge statistics of the labeling we have $p_{1,1}=0$, while $p_{0,1}=p_{1,0}=p_{1}$ is the density of the independent set that we usually denote by $\alpha$, and hence $p_{0,0}=1-2\alpha$.
Using the method outlined in Section 1.1, we will find a “Markovian independent set” with some density $\alpha$. Then one can design local algorithms that produce a larger independent set with density $\hat{\alpha}>\alpha$. In this paper we use a very basic algorithm. We call a vertex full-zero if its label is 0 and the labels of all its neighbors are also 0. Consider the induced subgraph over the full-zero vertices. We will see that the connected components are finite with probability $1$. Consequently, one may choose an independent set inside each finite component, and add these to our original independent set. It follows that the original Markovian independent set can be augmented by all one-vertex components and at least half of the remaining full-zero vertices. Since our starting independent set is Markovian, we can easily work out the density $\hat{\alpha}$ of our new independent set by computing the probability that a vertex is full-zero and the probability that it is an isolated full-zero vertex, see Section 4 for details.
One could, of course, consider more sophisticated algorithms to augment the Markovian independent set, resulting in a slightly higher density. However, the improvements would be minor, so we decided to stick with this basic method.
Table 2 shows the largest $\alpha$ that satisfies the second moment condition (3) in the context of independent sets and the corresponding $\hat{\alpha}$ gained by the basic improvement described above (for degrees in the range $3\leq d\leq 11$).
$d$ 3 4 5 6 7 8 9 10 11
$\alpha$ 0.41418 0.36580 0.33268 0.30776 0.28791 0.27151 0.25758 0.24552 0.23491
$\hat{\alpha}$ 0.42718 0.38339 0.35089 0.32525 0.30427 0.28663 0.27152 0.25837 0.24677
Table 2. Density of the Markovian and the augmented independent set
1.3. Further applications
To summarize, our approach is to use a stronger version of the second moment method to find some “typical” Markovian labeling, then perform certain local modifications. Note that the new (possibly non-Markovian) labeling may be a different “type of object” than the original Markovian labeling. In particular, we can start with the same Markovian independent set and create other objects, and hence proving the existence of more complex structures in random regular graphs. (For instance, suppose that we want to find large induced subgraphs with degree bound, say, $2$. It is not hard to design a local algorithm that produces such a subgraph from an independent set.)
In this paper we present one such application, concerning so-called star decompositions. In Section 5 we demonstrate how our approach can be used to prove that, with high probability, the edge set of a random $d$-regular graph can be decomposed into $k$-stars for certain pairs $d,k$.
Organization of the paper
Section 2 gathers the necessary facts about invariant and typical processes on the infinite tree. In Section 3 we examine the second moment condition for independent sets, while in Section 4 we discuss how the Markovian independent set can be augmented. Finally, Section 5 is concerned with star decompositions.
2. Typical processes
When studying labelings or colorings of random regular graphs, a useful tool is to study processes on the limiting graph, which is the $d$-regular infinite tree $T_{d}$. It is well known that $\mathbb{G}_{N,d}$ converges to $T_{d}$ locally (Benjamini–Schramm convergence) as $N\to\infty$. This is due to the fact that $\mathbb{G}_{N,d}$ is essentially large-girth with high probability: for every $\ell\in\mathbb{N}$ the number of cycles of length at most $\ell$ is $o(N)$.
2.1. Processes
By an invariant process on $T_{d}$ we mean a random coloring/labeling, the distribution of which is invariant under the automorphisms of $T_{d}$. The so-called tree-indexed Markov chains form a special class of invariant processes.
Definition 2.1.
Suppose that $\mathcal{X}$ is a finite set of colors/labels. A mapping $V(G)\to\mathcal{X}$ is a vertex labeling of the graph $G$. A random labeling $X\colon V(T_{d})\to\mathcal{X}$ is a $T_{d}$-indexed Markov chain if, conditioned on the label $X_{v}$ of a vertex $v\in V(T_{d})$, $X$ is conditionally independent on the $d$ subtrees obtained after deleting $v$.
The distribution $\mu$ of a $T_{d}$-indexed Markov chain is uniquely described by its marginal $\mu_{\mathrm{edge}}$ on two neighboring vertices of $T_{d}$. Note that $\mu_{\mathrm{edge}}\in\mathcal{P}_{\mathrm{edge}}^{\mathcal{X}}$, where we write $\mathcal{P}_{\mathrm{edge}}^{\mathcal{X}}$ for the set of distributions on $\mathcal{X}\times\mathcal{X}$ that are invariant under exchanging the two coordinates.
Definition 2.2.
An invariant process $\mu$ is said to be typical [^3] In [1, Def 1.1] the term strongly typical was used. if it can be mimicked on most $d$-regular graphs in the following sense. For every $\varepsilon>0$ and $R\in\mathbb{N}$ there exists $N_{0}=N_{0}(\varepsilon,R)$ such that for every $N>N_{0}$ the following holds for $\mathbb{G}_{N,d}$ with probability at least $1-\varepsilon$: the graph $\mathbb{G}_{N,d}$ has an $\mathcal{X}$-labeling with $R$-neighborhood statistics $\varepsilon$-close [^4] When we talk about the distance of discrete distributions, let us use, say, the total variation distance. to that of $\mu$.
2.2. Entropy
Given an exchangeable distribution $\pi\in\mathcal{P}_{\mathrm{edge}}^{\mathcal{X}}$ on $\mathcal{X}\times\mathcal{X}$, we write $\pi_{\mathrm{vtx}}$ for its marginal on any of the two coordinates, and define
(4) $\Sigma(\pi)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\frac{d}{2}H(\pi)-(d-1)H(\pi_{\mathrm{vtx}}).$
See (2) for how Shannon entropy $H(\cdot)$ is defined. The motivation for this definition is that $\Sigma(\pi)$ gives the exponential rate of the expected number of $\mathcal{X}$-labelings of $\mathbb{G}^{\textrm{conf}}_{N,d}$ with edge statistics close to $\pi$. In particular, if $\Sigma(\pi)<0$, then $\mathbb{G}^{\textrm{conf}}_{N,d}$ may have such a labeling only with exponentially small probability (in $N$). This fact is what we usually refer to as the first moment method in random regular graphs.
When we wish to compute (the exponential rate of) the second moment of the number of such labelings, then we need to determine the expected number of pairs of labelings (one red and one blue), and hence consider $\Sigma(\nu)$ for distributions $\nu$ over
[…]
We will use the following stronger form of the second moment method.
Lemma 2.3 (Backhausz–Bordenave–Szegedy, Thm 1.15 of [1]).
Suppose that $\mu$ is a tree-indexed Markov chain on $T_{d}$ with the property that for any coupling $\nu\in\mathcal{P}_{\mathrm{edge}}^{\mathcal{X}\times\mathcal{X}}$ of $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mu_{\mathrm{edge}}$} and $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mu_{\mathrm{edge}}$} we have
(5) $\Sigma(\nu)\leq 2\Sigma(\mu_{\mathrm{edge}}).$
Then $\mu$ is (strongly) typical.
Remark 2.4.
[…]
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}i}$ $:\quad\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}red} label of the {first} vertex;}$
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}j}$ $:\quad\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}blue} label of the {first} vertex;}$
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}k}$ $:\quad\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}red} label of the {second} vertex;}$
[…]
We need to consider distributions $\nu$ over $\mathcal{X}^{4}$ that is invariant under exchanging the first and second vertex [^5] Note that $\nu$ does not have to be invariant under exchanging the two colors. and its marginal for any color is $\mu_{\mathrm{edge}}$. We will use the notation
$p_{ij,k\ell}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\nu\big(\{(i,j,k,\ell)\}\big).$
Note that the comma separates the labels of the first and second vertex of the edge. With this notation, the property that $\nu$ is exchangeable can be expressed as
(6) $p_{ij,k\ell}=p_{k\ell,ij}\quad(\forall i,j,k,\ell\in\mathcal{X}).$
As for the condition on the marginals, we have
(7) $\sum_{j,\ell\in\mathcal{X}}p_{ij,k\ell}$ $=p_{i,k}\quad(\forall i,k\in\mathcal{X});$
(8) $\sum_{i,k\in\mathcal{X}}p_{ij,k\ell}$ $=p_{j,\ell}\quad(\forall j,\ell\in\mathcal{X}).$
Remark 2.5.
Consider the independent coupling $\nu^{\star}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mu_{\mathrm{edge}}\times\mu_{\mathrm{edge}}$ for which $p_{ij,k\ell}=p_{i,k}p_{j,\ell}$ (that is, the red and blue labels are independent). We clearly have
$\Sigma(\nu^{\star})=2\Sigma(\mu_{\mathrm{edge}}).$
Therefore, the second moment condition (5) can also be written as $\Sigma(\nu)\leq\Sigma(\nu^{\star}),\forall\nu$.
Note that if we choose $\nu$ to be the distribution of two identical copies of $\mu_{\mathrm{edge}}$ (that is, the blue and red labels are always the same), then $\Sigma(\nu)=\Sigma(\mu_{\mathrm{edge}})$. Therefore, if the second moment condition holds, then we necessarily have the “first moment condition” $\Sigma(\mu_{\mathrm{edge}})\geq 0$.
2.3. Factors
As we mentioned in the introduction, many previous lower bounds on $\alpha^{\star}_{d}$ used (randomized) local algorithms to construct independent sets. These algorithms can be applied to any graphs of essentially large girth, not only to random regular graphs. In the language of graph limits, such local algorithms are equivalent to the so-called FIID (factor of IID) processes. See [2] for the precise definition and some basic facts about these processes.
It is easy to see that FIID processes are (strongly) typical. In fact, one can start from a typical process and apply a factor mapping; the resulting process is always typical. More specifically, suppose that $X\colon V(T_{d})\to\mathcal{X}$ is an invariant process on $T_{d}$, and let $Z\colon V(T_{d})\to[0,1]$ be an IID process (also independent from $X$). In this context, by a factor mapping we mean an $\operatorname{Aut}(T_{d})$-equivariant measurable mapping $\Psi\colon\big(\mathcal{X}\times[0,1]\big)^{V(T_{d})}\to\mathcal{Y}^{V(T_{d})}$. Clearly, $Y\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\Psi(X,Z)\colon V(T_{d})\to\mathcal{Y}$ is also invariant under $\operatorname{Aut}(T_{d})$. Moreover, it is an easy fact that if $X$ is typical, then so is $Y$.
Essentially, it suffices to consider block factors, where the new label of a vertex $v$ is determined by the original labels on $B_{R}(v)$. In essence, $Y_{v}$ is computed based on the original labels $X_{u}$ and the IID labels $Z_{u}$ for $u\in B_{R}(v)$, and each vertex has to use the same measurable “rule”.
In our setting, we will start from some typical Markovian labeling $X_{v}$, and obtain $Y_{v}$ using a fixed local algorithm based on the $X$-labels around $v$, possibly using some independent random choices along the way. The resulting labeling is, therefore, also typical.
3. Second moment for independent sets
In this section we apply the machinery presented in Section 2 to independent sets.
Definition 3.1.
We define an invariant independent set as an invariant process on $T_{d}$ such that each label is 0 or $1$ with no adjacent $1$’s. If, in addition, the process has the Markov property (see Definition 2.1), then we call it a Markovian independent set.
By the density we mean the probability that a given vertex has label $1$. Note that a Markovian independent set is uniquely determined by its density.
Fix some degree $d\geq 3$ and let $\mu=\mu(d,\alpha)$ denote the distribution of the Markovian independent set on $T_{d}$ with density $\alpha$. Our goal in this section is to prove that $\mu$ is typical for some $\alpha$ (as large as possible).
The vertex and edge marginals of $\mu$ are as follows:
$\mu_{\mathrm{vtx}}$ $:\quad p_{1}=\alpha;\;p_{0}=1-\alpha;$
$\mu_{\mathrm{edge}}$ $:\quad p_{1,1}=0;\;p_{0,1}=p_{1,0}=\alpha;\;p_{0,0}=1-2\alpha.$
Now let $\nu$ be an arbitrary coupling of $\mu_{\mathrm{edge}}$ and $\mu_{\mathrm{edge}}$. (One may think of two random independent sets, one red and one blue, both of density $\alpha$.) As explained in Section 2, we need to consider distributions $\nu$ over $\{0,1\}^{4}$ for which the probabilities $p_{ij,k\ell}$ satisfy (6), (7), (8).
Let $\beta\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=p_{11}$. (This corresponds to the density of the intersection of the blue and red independent sets. Note that this is different from $p_{1,1}=0$, which is the probability that two neighboring vertices have 0 blue labels.) Then we have
$\nu_{\mathrm{vtx}}:\quad p_{11}=\beta;\;p_{01}=p_{10}=\alpha-\beta;\;p_{00}=1-2\alpha+\beta.$
It follows that
$p_{00,11}=p_{11,00}=\beta.$
Now let $\gamma\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=p_{10,01}=p_{01,10}$. Then
$p_{01,00}+p_{01,10}=p_{01}=\alpha-\beta\text{, and hence }p_{01,00}=\alpha-\beta-\gamma=p_{00,01}.$
Similarly,
$p_{10,00}+p_{10,01}=p_{10}=\alpha-\beta\text{, and hence }p_{10,00}=\alpha-\beta-\gamma=p_{00,10}.$
Then the remaining probability is $p_{00,00}=1-4\alpha+2\beta+2\gamma$, and we get that
$H(\nu_{\mathrm{vtx}})$ $=h(\beta)+2h(\alpha-\beta)+h(1-2\alpha+\beta);$
$H(\nu)$ $=2h(\beta)+2h(\gamma)+4h(\alpha-\beta-\gamma)+h(1-4\alpha+2\beta+2\gamma).$
For fixed $\alpha,\beta$, the function $\gamma\mapsto H(\nu)$ has the following derivative:
$2\log\gamma+4\log(\alpha-\beta-\gamma)-2\log(1-4\alpha+2\beta+2\gamma).$
It is easy to see that this is is monotone decreasing in $\gamma\in[0,\alpha-\beta]$ and it vanishes if and only if
$(\alpha-\beta-\gamma)^{2}=\gamma(1-4\alpha+2\beta+2\gamma),$
which leads to
(9) $\gamma=(\alpha-1/2)+\sqrt{(\alpha-1/2)^{2}+(\alpha-\beta)^{2}}.$
Therefore, $H(\nu)$ attains its maximum at (9) for any fixed $\alpha,\beta$.
In order to check the second moment condition of Lemma 2.3, we need to maximize
$\Sigma(\nu)=\frac{d}{2}H(\nu)-(d-1)H(\nu_{\mathrm{vtx}})$
and show that the maximum is attained at the independent coupling $\nu^{\star}$ (see Remark 2.5). Since $H(\nu_{\mathrm{vtx}})$ does not depend on $\gamma$, we may make the substitution (9) for $\gamma$. We get a function in $\alpha,\beta$. In fact, if we treat $d$ and $\alpha$ as parameters and $\beta$ as the single variable, we get the following function $f_{d,\alpha}(\beta)$ expressing $\max_{\gamma}\Sigma(\nu)$:
[…]
The second moment condition is satisfied for a given $\alpha$ if this function attains its maximum at $\beta=\alpha^{2}$ (corresponding to the independent coupling). We summarize our findings in the following theorem.
Theorem 3.2.
Let $d\geq 3$ be an integer and let $\alpha\in(0,1/2)$. Suppose that the function $f_{d,\alpha}\colon[0,\alpha]\to\mathbb{R}$ attains its maximum at $\alpha^{2}$. Then the Markovian independent set of density $\alpha$ is a typical process on $T_{d}$.
It turns out that, when $\alpha$ is in the relevant region, the function $f_{d,\alpha}$ has two local maxima, one at $\alpha^{2}$ and another one above $\alpha^{2}$. (It is actually close to $\alpha$ for large $d$). So one needs to check which local maximum is larger, and find (the largest possible) $\alpha$ for which the global maximum is at $\beta=\alpha^{2}$ (corresponding to the independent coupling). This is quite straightforward to do with the help of a computer.
See Table 2 in the introduction for the largest $\alpha$ satisfying the second moment condition for small degrees.
4. Augmented independent sets
In this section we discuss the local improvements that one can make to an independent set with the spatial Markov property. Let us start from a Markovian independent set of density $\alpha$ (see Definition 3.1). If we have two neighboring vertices, then the conditional probability that one has label 0 conditioned on the other having label 0 is equal to
$p\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\frac{1-2\alpha}{1-\alpha}.$
In the introduction we defined a full-zero vertex as a vertex whose label, as well as all its neighbors’ labels, are 0. Clearly, the probability that a vertex is full-zero is
(11) $(1-\alpha)p^{d}.$
Furthermore, the probability that a vertex is an isolated full-zero vertex (that is, a full-zero vertex with no full-zero neighbor) is
(12) $(1-\alpha)p^{d}\big(1-p^{d-1}\big)^{d}.$
Now consider the induced subgraph over all full-zero vertices. We will need that the connected components of this subgraph are finite (almost surely). Note that the conditional probability that a given neighbor of a full-zero vertex is also full-zero is equal to $p^{d-1}$. Therefore, if the condition
$p^{d-1}\leq\frac{1}{d-1}$
holds, then the corresponding Galton–Watson process dies out almost surely, and hence we have finite components. This condition is always comfortably satisfied whenever $\alpha$ satisfies the second moment condition. Therefore, we may assume that the components are finite.
It is easy to design a randomized local algorithm that selects an independent set inside each finite component with size at least half of the component. [^6] The components are trees and hence bipartite. Take a “canonical” proper 2-coloring and choose the larger color class. If they have the same size, choose the one in which the sum of IID labels is larger. Let us add these to the original independent set. Of course, the resulting set remains to be independent. What can we say about its density? We augmented the original set with all the one-vertex components (i.e., isolated full-zero vertices) and at least half of the remaining full-zero vertices. Therefore, using (11) and (12), the new density is at least
(13) $\hat{\alpha}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\alpha+(1-\alpha)p^{d}\frac{1+\big(1-p^{d-1}\big)^{d}}{2}.$
We summarize our findings in the next theorem.
Theorem 4.1.
Let $d\geq 3$ be an integer and $\alpha\in(0,1/2)$, and set $p=(1-2\alpha)/(1-\alpha)$. Suppose that the function $f_{d,\alpha}\colon[0,\alpha]\to\mathbb{R}$ attains its maximum at $\alpha^{2}$, and $(d-1)p^{d-1}\leq 1$. Then there exists a typical independent set on $T_{d}$ with density $\hat{\alpha}$ as in (13).
See Table 2 in the introduction for the density $\hat{\alpha}$ of the augmented independent set for small degrees.
One could get a slightly higher density by also computing the probability of components with three vertices, or by considering further local improvements but they would only lead to marginal gains.
5. Star decompositions
To further demonstrate the strength of our approach, in this section we consider the following problem. Given an integer $d/2<k<d$, can the edges of $\mathbb{G}_{N,d}$ be partitioned into edge-disjoint stars, each containing $k$ edges? We restrict ourselves to those $N$ for which the number of edges ($Nd/2$) is divisible by $k$. If such a partition exists with probability $1-o_{N}(1)$, then we say that $\mathbb{G}_{N,d}$ asymptotically almost surely (a.a.s.) has a $k$-star decomposition.
The case $d=4,k=3$ was settled (affirmatively) in [7]. Later, that method was extended in [6], and a sufficient condition was given for general pairs $d,k$. Both papers used a second moment method and checking the required conditions lead to very technical computations.
The problem is closely related to the asymptotic independence ratio $\alpha^{\star}_{d}$ of $\mathbb{G}_{N,d}$. More specifically, if $\mathbb{G}_{N,d}$ a.a.s. has a $k$-star decomposition, then
$\alpha^{\star}_{d}\geq 1-\frac{d}{2k}\text{, or equivalently: }k\leq\frac{d}{2(1-\alpha^{\star}_{d})}.$
This gives a necessary condition for the existence of a $k$-star decomposition. It was conjectured in [6] that this is essentially the only requirement. This was confirmed for most degrees (in an asymptotic sense) in [13]. In that paper we showed that the existence of large independent sets with an additional property (called thinness) actually implies the existence of star decompositions provided that the graph has some properties regarding the edge density of its induced subgraphs. We will use this general result here. First we define the notion of thin sets in graphs.
Definition 5.1.
Given a positive integer $\hat{d}<d$, we say that a set $A\subset V(G)$ is $\hat{d}-$thin in $G$ if every $v\in V(G)\setminus A$ has at most $\hat{d}$ neighbors in $A$.
Then we have the following sufficient condition for the existence of a $k$-star decomposition.
Lemma 5.2 ([13]).
Let $G$ be a $d$-regular graph on $N$ vertices and let $k>d/2$. Set
$\alpha_{d,k}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=1-\frac{d}{2k}.$
For $U\subseteq V(G)$ let $e[U]$ denote the number of edges of $G$ inside $U$. Suppose that $A\subset V(G)$ satisfies the following conditions for a positive integer $\hat{d}<k$ and a real number $0<c<1$.
  1. (i)
    AA is an independent set of density $\alpha_{d,k}$.
  2. (ii)
    AA is $\hat{d}-$thin.
  3. (iii)
    For any set $U\subseteq A^{\mathrm{c}}=V(G)\setminus A$ with $|U|\leq cN$ we have
    $e[U]\leq(d-k)|U|.$
  4. (iv)
    For any set $W\subseteq A^{\mathrm{c}}$ with $|W|<(1-\alpha_{d,k}-c)N$ we have
    $e[W]\leq(k-\hat{d})|W|.$
Then $G$ has a $k$-star decomposition.
If we choose the parameters properly, then conditions (iii) and (iv) actually hold true a.a.s. for random regular graphs. In fact, one can simply check the numerical condition of [13, Cor 3.4] to check (iii) and (iv) in specific cases. We omit the details here as this is not the main focus of the current paper. Instead, we focus on the first two conditions.
To find a set $A$ satisfying (i) and (ii), we start from a Markovian independent set of density $\alpha$. As before, $p\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(1-2\alpha)/(1-\alpha)$ is the conditional probability that a vertex has label 0 conditioned on the event that a fixed neighbor has label 0. Then, by the Markov property, the probability that a vertex with label 0 has $\ell$ neighbors with label $1$ is the following:
$\binom{d}{\ell}(1-p)^{\ell}p^{d-\ell}.$
For any vertex $v\notin A$ with $\ell>\hat{d}$ neighbors in $A$, we may remove $\ell-\hat{d}$ vertices from $A$ to ensure that $v$ does not violate the thinness condition. (Note that, although we created some new outside vertices during this removal process, they all have 0 neighbors in $A$.) This simple local algorithm removes a subset of $A$ of density at most
$(1-\alpha)\sum_{\ell=\hat{d}+1}^{d}(\ell-\hat{d})\binom{d}{\ell}(1-p)^{\ell}p^{d-\ell}.$
Therefore, we have proved the following.
Lemma 5.3.
Assume that the Markovian independent set of density $\alpha$ is typical, and set $p\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(1-2\alpha)/(1-\alpha)$. For $\hat{d}<d$ there exists a $\hat{d}-$thin invariant independent set $A_{\mathrm{thin}}\subseteq A$ of density at least
$\alpha_{\mathrm{thin}}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\alpha-(1-\alpha)\sum_{\ell=\hat{d}+1}^{d}(\ell-\hat{d})\binom{d}{\ell}(1-p)^{\ell}p^{d-\ell},$
which is also typical.
For a specific example, consider the degree $d=33$. In this case, the density of the Markovian independent set is $\alpha\approx 0.13195395$, while the corresponding conditional probability is $p\approx 0.84798738$. Then, by Lemma 5.3 we get the following densities for $\hat{d}-$thin independent sets for $\hat{d}=8,9,\ldots,14$:
$\hat{d}$ deleted $\alpha_{\mathrm{thin}}$ star decomposition
88 $0.073901$ $0.058052$ implies $k=17$ with $c=0.1$
99 $0.027726$ $0.104227$ implies $k=18$ with $c=0.1$
$10$ $0.009286$ $0.122667$
$11$ $0.002778$ $0.129175$
$12$ $0.000743$ $0.131210$
$13$ $0.000177$ $0.131776$ implies $k=19$ with $c=0.1$
$14$ $0.000038$ $0.131915$
Then one can use Lemma 5.2 to prove the existence of $k$-star decompositions in $\mathbb{G}_{N,33}$ for $k=17,18,19$. (Note that star decompositions cannot exist for $k\geq 20$ because $\alpha^{\star}_{33}\leq 0.14401<0.175=1-\frac{33}{40}$.) As we mentioned, conditions (iii) and (iv) can be easily checked (for the choice $c=0.1$) with the tools in [13]. This demonstrates that, using this method, one can reproduce many results proven in [6] with substantially smaller technical difficulty.